# Why Foundation Models in Pathology Are Failing

### v1

@@ 2. Empirical Evidence of Weaknesses
Lack of Robustness and Site Bias – De Jong et al. [7] systematically evaluated the robustness and generalization of ten leading pathology foundation models across multiple institutions and datasets using a newly defined Robustness Index (RI), which quantifies whether model embeddings cluster more strongly by biological class or by medical center. The RI compares within-class versus within-center similarity (with RI>>1 indicating true biological robustness). Among all tested models, only Virchow2 achieved RI $\approx$ 1.2—meaning biological structure dominated site-specific bias—whereas all others had RI ¡ 1 (e.g., UNI $\approx$ 0.9, Phikon-v2 $\approx$ 0.7). Embeddings from most models, therefore, grouped primarily by hospital or scanner rather than by cancer type, leading to large performance drops on unseen centers. The study concludes that current pathology foundation models remain fragile, confounded, and insufficiently domain-robust, underscoring the urgent need for cross-institutional validation and bias-resilient architectures before clinical use.
@@
Security and Safety Vulnerabilities – Wang et al. [12] introduced Universal and Transferable Adversarial Perturbations (UTAP), imperceptible noise patterns that collapse FM embeddings across architectures. These universal attacks threaten clinical reliability. In histopathology, these perturbations have a dual interpretation go beyond malicious attacks that are clearly a security risk. They have a real-world noise analogue; they approximate the small, systematic variations that arise naturally in the imaging pipeline:
@@ 3. Why Are Pathology FMs Failing?
Vision Transformers (ViTs) split images into small patches (e.g., 16×16) to convert them into manageable token sequences for self-attention. This is computationally efficient but semantically costly: the model initially loses the global spatial structure (paramount for tissue morphology) and must learn it back from data. In pathology, where diagnostic meaning resides in multimagnification architecture, this design leads to models that see textures but not tissues.

### v2

@@ 2. Empirical Evidence of Weaknesses
Lack of Robustness and Site Bias – De Jong et al. [7] systematically evaluated the robustness and generalization of ten leading pathology foundation models across multiple institutions and datasets using a newly defined Robustness Index (RI), which quantifies whether model embeddings cluster more strongly by biological class or by medical center. The RI compares within-class versus within-center similarity (with RI>>1 indicating true biological robustness). Among all tested models, only Virchow2 achieved RI $\approx$ 1.2—meaning biological structure dominated site-specific bias—whereas all others had RI << 1 (e.g., UNI $\approx$ 0.9, Phikon-v2 $\approx$ 0.7). Embeddings from most models, therefore, grouped primarily by hospital or scanner rather than by cancer type, leading to large performance drops on unseen centers. The study concludes that current pathology foundation models remain fragile, confounded, and insufficiently domain-robust, underscoring the urgent need for cross-institutional validation and bias-resilient architectures before clinical use.
@@
Security and Safety Vulnerabilities – Wang et al. [12] introduced Universal and Transferable Adversarial Perturbations (UTAP), imperceptible noise patterns that collapse FM embeddings across architectures. These universal attacks threaten clinical reliability. The paper shows that visually imperceptible perturbations at modest pixel bounds can severely degrade representation quality, collapsing accuracy from $\approx\!97$% to $\approx 12$% on the attacked model and transferring across black-box models (e.g., $96.42\%\rightarrow 48.69\%$, $97.23\%\rightarrow 25.52\%$), while also harming out-of-distribution datasets. These attacks break the learned feature manifold rather than merely flipping labels, are easy to learn ($\approx 900$ patches, a few minutes on an RTX,4090), and generalize widely. Thus, the high clean-test accuracy of ViT-based patch encoders can be deceptive, leaving deployed pathology systems vulnerable and posing significant safety risks to patients.
In histopathology, these perturbations pose significant safety risks and have a dual interpretation that go beyond malicious attacks that are clearly a security risk. They have a real-world noise analogue; they approximate the small, systematic variations that arise naturally in the imaging pipeline:
@@ 5. Ethical Considerations
While empirical data show clear evidence of publication bias (i.e., fewer studies with negative or null results appear in the literature) [18], there is currently no robust published study demonstrating that journals routinely invite non-expert or sympathetic reviewers to favour a particular outcome. Nevertheless, the possibility remains plausible given structural incentives in AI/ML publishing and merits caution.
@@
Several meta-studies in machine-learning and medical-imaging have documented incomplete method disclosure (e.g., hyper-parameters, seeds, data-splits) and pervasive data-leakage biases (e.g., subject-wise leakage, tile reuse) that systematically inflate reported model performance [19, 20, 21, 22, 23].
@@
Recent surveys of healthcare executives show strong optimism and investment in AI — for example, more than two-thirds of C-suite leaders report their organizations invest in AI, yet only  12-13% believe current algorithms are robust enough or have a clear strategy for clinical integration [24, 25, 26].